Red light with a wavelength of 725 nm shines through a single slit, and forms a pattern on a screen placed 1.25 m away. If the width of the slit is 18.5 micrometers, find the width of the central maximum (a) in degrees and (b) in centimeters
1
Expert's answer
2019-04-18T10:51:04-0400

The angular position "\theta_m" of the first minimum is determined by the equation "\sin \theta_m = \lambda / d", where "\lambda = 725 \times 10^{-9}\, \text{m}" is the wavelength, and "d = 18.5 \times 10^{-6}\, \text{m}" is the width of the slit. We obtain "\sin \theta_m = 0.725 / 18.5 \approx 0.0392". In view of the smallness of this number, we have "\sin \theta_m \approx \theta_m \approx 0.0392". The angular width of the central maximum is twice this angle: "\theta = 2 \theta_m \approx 0.0784". The angular width in degrees is then "360\, \theta / 2 \pi \approx 4.5^\circ". The width in centimetres is proportional to the distance to the screen "R = 125\, \text{cm}" and is given approximately by "\theta R = 0.0784 \times 125\, \text{cm} = 9.8\, \text{cm}".


Answer: (a) "4.5^\circ", (b) "9.8\, \text{cm}".